OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 5i

OpenStudy (cruffo):

Since 2 + 5i is a zero, then 2 - 5i is also a zero. Thus, since there are four zeros, the minimum degree will be 4, that is, the polynomial will be of the form \(P(x) = ax^4 + bx^3 + cx^2 +dx + e\)

OpenStudy (anonymous):

ok...but i cant get any of the answers that are the choices..:(

OpenStudy (cruffo):

Yah, there is still a bit of work to do... You have to get the factors of the polynomial from the zeros, then multiply all the factors together to get the polynomial.

OpenStudy (cruffo):

Now, to get the factors from the zeros : Since x = 4 is a zero, then x - 4 = 0 so (x - 4) is a factor. Since x = -8 is a zero, then x + 8 = 0, so (x + 8) is a factor. Since x = 2 +5i is a zero, then x - 2 - 5i = 0, so (x - 2 - 5i) is a factor. Since x = 2 +5i is a zero, then x - 2 - 5i = 0, so (x - 2 - 5i) is a factor. Since x = 2 - 5i is a zero, then x - 2 + 5i = 0, so (x - 2 + 5i) is a factor. So the polynomial can be written as \[P(x) = (x - 4)(x - 8)(x-2-5i)(x - 2 + 5i)\]

OpenStudy (anonymous):

f(x) = x4 - 6x3 - 20x2 + 122x - 928 f(x) = x4 - 19x2 + 244x - 928 f(x) = x4 - 6x3 + 20x2 - 122x + 928 f(x) = x4 - 61x2 + 244x - 928

OpenStudy (anonymous):

those are the choices

OpenStudy (cruffo):

Now, to multiply the factors (we can do this in pairs): \[(x-4)(x+8) = x^2 + 4x - 32\] and \[(x-2-5i)(x-2+5i) = x^2-4x +25\] So \[P(x) = (x^2+4x-32)(x^2-4x+25)\]

OpenStudy (cruffo):

To finish, just multiply these two trinomials, gather liker terms, and you should get one of the answer choices.

OpenStudy (cruffo):

I'm not getting any of the answer choices...

OpenStudy (anonymous):

x4 - 4x3 + 25x2 + 4x3 - 16x2 + 100x - 32x2 + 128x - 800 x4 - 23x2 + 228x -800 is what i get

OpenStudy (anonymous):

what did you get? idk how this is getting so wrong

OpenStudy (cruffo):

I made a mistake when I multiplied \((x-2-5i)(x-2+5i)\). I should have gotten \(x^2 - 4x + 29\). http://www.wolframalpha.com/input/?i=factor+++x%5E4+-+19x%5E2+%2B+244x+-+928

OpenStudy (anonymous):

so its x^4 - 19x^2 + 244x - 928

OpenStudy (anonymous):

???

OpenStudy (cruffo):

Yes. That is correct. Sorry for the delay.